OpenStudy (anonymous):

2x + 4y = 12 2x + y = 3 2x + 3y = 8 -x + y = -4 4x + 2y = 12 4x + 2y = 4 2x + 3y = 4 6x + 9y = 12 How do i solve them

OpenStudy (anonymous):

by the simultaneous way look i'll solve u one as an example & u do the others on the same Pattern...........so let's take no.2 2x+3y=8 in the simultaneous way the value of (X) or (y) -x+y=-4 must be equal in the 2 equations to be cancelled therefore we will multiply equation 2 time 2 (-x+y=-4) X 2 = -2x+2y=-8 so......2x+3y=8 -2x+2y=-8 so by adding the 2 equations we get that 5y=0 so y=0 so x=4

OpenStudy (anonymous):

thankz

OpenStudy (anonymous):

wlcm :)